A piston-cylinder arrangement contains a gas at P1 = 600 KPa, V1 = 0.10m3 and expand to V2 = 0.5 m3. Compute the work done by the gas if the process follows the relation (a) pV1.4 = c, and (b) p = -300V + 630, where p is in KPa and V in m3.


1
Expert's answer
2021-01-21T03:52:36-0500

a)

PV1.4 =c

V =1.4

C= p1V1r

=600 *(0.1)1.4

C=23.886

work done w = c(V21-r-V11.r)/(1-r)

=23.886(0.5(1-1.4)-0.10(1-1.4))/1-1.4

answer = 71.203 kg



b)p- -300V + 630

w="\int"pdv

="\int"( -300V + 630)dv.................. from 0.1 to 0.5

=[(-300V2/2)/2 +630V]................. from 0.1 to 0.5

=[(-300/2)(0.52-0.12)/2 +630(0.5-0.1)]

answer w=216 kg